the sum of 2 numbers is 15 less than twice the first number. their difference is 5 less than twice the second number. find each of the numbers
Answer by RAY100(1637)

(1) a+b = 2a -15,,,,,,-a+b =-15
.
(2) a-b =2b-5,,,,,,,,,,,,,a-3b = -5
.
add(1) + (2),,,,,,,,,,,,,,,,,,-2b = -20
.
b= 10
.
substitute in (1) -a +(10) = -15
.
-a = -25
.
a =25
.
check (1) 25 +10 =2*25 -15 =35 ,,,ok
.
(2) 25 -10 = 2*10 -5 = 15,,,,,ok